Kane & Associates is a local legal entity employing six solicitors, four paralegals and six administrative staff. The entity has a costing system that assigns costs to clients. There are two departments — legal support and commercial. The entity uses a different indirect cost rate for each department. The following information was taken from the budget for the coming year.

Legal support
department   

Commercial

department

Planned solicitor hours 3,200 12,400
Planned paralegal hours       4,100
Number of clients 300 950
Legal supplies       $16,200 $27,000
Salaries — solicitors and paralegals       $162,000 $810,000
Indirect costs $108,000 $450,000

  

The indirect costs are allocated to the clients as follows.

●       Departmental indirect costs — number of solicitor and paralegal hours.  
●       Legal supplies are charged on a per client basis.  

The entity uses a cost-plus pricing policy with a 20 per cent mark-up.

Client Case Number FR110 has been completed and is ready for the invoice to be sent to the client. The following costs and times were recorded for the client.

●       Solicitor hours — Support 65 hours.  
●       Solicitor hours — Commercial 187 hours.  
●       Direct cost of solicitors — $14,000.  

A/ Calculate the indirect cost rate for each department. (Round answers to 2 decimal places, e.g. 15.25.)
B/ calculate the indirect cost rate for legal supplies.
C/ determine the total cost for client case FR110.
D/ calculate the amount the client should be invoiced for the work undertaking.

this my second time posting this question , please take your time to answer these questions correctly.

thank you
Answer #1

A) Indirect cost rate

Legal support department:-

As it is provided that departmental indirect cost is provided on the basis of number of solicitor and legal hours

Indirect cost rate = total indirect cost/ solicitor and legal hours

= 108000/3200

= 33.75 per hour

Commercial department = total indirect cost / solicitor and legal hours

= 450000/(12400+4100)

= 27.27 per hour

B) Legal supplies rate

Legal support department:-

As the same is on a per client basis.

Indirect cost rate = legal supplies/ number of clients

= 16200/ 300

= 54 per client

Commercial department

Indirect cost rate = legal supplies/ number of clients

= 27000/950

= 28.42 per client

C) Total cost for client case FR110 = indirect cost of legal support department + indirect cost of commercial department + direct cost provided + legal supplies cost

= (65*33.75)+(187*27.27)+54+28.42+14000

=2193.75+5100+54+28.42+14000

=$21376.17

D) invoice amount for client = cost + 20% mark mark up

= 213676.17+(213676.17*20%)

= 25651.40